PUMP-PROBE PHOTOIONIZATION AND MASS SPECTROSCOPY OF PENTAMETHYLCYCLOPENTA-DIENE

\underline{F. RUDAKOV}, P. WEBER, Department of Chemistry, Brown University, Providence, RI 02912.

The ultrafast radiationless decays in cyclic dienes normally proceed through conical intersections of potential energy surfaces. Time-resolved experiments on pentamethylcyclopentadiene were performed by exciting the first excited state with femtosecond pulses at 267 nm. Photoionization with a time-delayed probe pulse yields delay-time dependent mass and photoelectron spectra that reveal the ultrafast character of the curve crossing dynamics.